jack is 4 years older than twice his sister's age.in 4 years time he will be 11 years older than her. how old is joseph now?
Answer by stanbon(75887)

jack is 4 years older than twice his sister's age.
in 4 years time he will be 11 years older than her.
how old is joseph now?
----
Equations:
j = 2s+4
(j+4) = (s+4)+11
-----------------------
Substitute for "j" and solve for "s":
(2s+4+4) = s+15
2s + 8 = s+15
s = 7 yrs (Sister's age now)
-----
Since j = 2s+4, j = 16+4 = 20 yrs (Jack's age now)
========================
Cheers,
Stan H.